If x+y+z=90°,Prove that 1-sin^2x-sin^2y-sin^2z-2sinx siny sinz=0

Grade:12th pass

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Please consider all angles equal 30, 30, 30

LHS = 1- 3(1/8) - 2(1/8)= 1-5/8 = 3/8 != RHS

Please check and upload correct question

Regards